# Plasmonic metamaterial time crystal

### v1

@@ Abstract
Periodically driven optical materials and metamaterials have recently emerged as a promising platform for realizing photonic time crystals (PTCs)—systems whose optical properties are strongly and periodically modulated on time scales comparable to the optical cycle of light Asgari_2024 . These time-varying structures are the temporal counterparts of spatial photonic crystals (SPCs), for which a large and periodic dielectric contrast is achieved spatially on wavelength scales. Just as SPCs have revolutionized control over light–matter interactions by engineering the photonic density of states in space Yablonovitch_1987 ; John_1987 ; Joannopoulos_2011 ; Lu_2014 , PTCs promise comparable breakthroughs from a fundamentally new perspective: a temporal one Lustig_2018 ; Lyubarov_2022 ; Dikopoltsev_2022 ; Li_2023 ; Wang_2024 ; Park_2025 . However, harnessing such phenomena at optical frequencies poses severe experimental challenges Asgari_2024 ; Hayran_2022 ; Saha_2023 ; Hayran_2025 , as it requires order-unity modulation depths of the optical properties at optical cycle rates, a regime that has remained elusive to date.
Here, we report the first optical realization of a photonic time crystal, achieved with a surface plasmon cavity metamaterial operating at Terahertz frequencies. We demonstrate strong (near-unity) and coherent (sub-optical cycle) periodic driving of the plasmonic metamaterial enabled by field-induced dynamical modulation of the carriers’ kinetic energy and effective mass—reaching up to 80% of their rest mass, an exceptionally high value that forms the basis for time crystalline phenomena with plasmons Kiselev_2024 ; Feinberg_2025 . Our experimentally informed theory reveals rich physics within the experimentally accessible parameter regime of this system, including parametric amplification and entangled plasmon generation, and establishes a robust new platform for time-domain photonics.
@@ Introduction
Systems subjected to periodic temporal modulations, known as Floquet systems, have been extensively studied across diverse platforms, including mechanical Trainiti_2016 , atomic Eckardt_2017 , and solid-state Wang_2013 ; Nova_2016 ; Cavalleri_2017 ; McIver_2019 ; Oka_2019 ; Rudner_2020 . In such systems, periodic driving is known to provide fascinating properties not accessible in equilibrium Bukov_2015 and the ability to tailor emergent properties through Floquet engineering of effective Hamiltonians Goldman_2014 ; Weitenberg_2021 . Over the past few years, these ideas have increasingly influenced the area of photonics. Periodic temporal modulation of optical systems enables time, in addition to space, to be utilized as a synthetic dimension for the tailoring of optical eigenmodes and response functions, an approach described within the framework of four-dimensional or Floquet metamaterials Engheta_2020 ; Yin_2022 ; Galiffi_2022 ; Sapienza_2023 ; Engheta_2023 ; Horsley_2023 ; Sisler_2024 . For spatially homogeneous yet ultrafast temporal modulations approaching the timescale of the optical cycle, the related concept of photonic time crystals (PTCs) has recently emerged Asgari_2024 ; Lustig_2018 ; Lyubarov_2022 ; Dikopoltsev_2022 ; Li_2023 ; Wang_2024 ; Park_2025 ; Kiselev_2024 ; Feinberg_2025 ; Pan_2023 ; Hayran_2024 ; Khurgin_2024 . As depicted in Figure 1a, PTCs can be viewed as the temporal analogues of spatial photonic crystals (SPCs) Yablonovitch_1987 ; John_1987 ; Joannopoulos_2011 ; Lu_2014 by permuting space and time coordinates on one hand and momentum and frequency on the other. In SPCs, spatial periodicity generates band replicas along the momentum axis and opens frequency gaps ($\omega$-gaps), whereas in PTCs, temporal periodicity produces replicas along the frequency axis and opens momentum gaps ($k$-gaps). Wave-behavior within or in the vicinity of these gaps is key in governing light–matter interactions and differ fundamentally between the two cases. Inside the gaps, wave solutions reduce to exponentially decaying or growing modes — spatially in SPCs and temporally in PTCs (Fig. 1a). In SPCs, only the decaying solution is physically allowed, a consequence of their time-invariant and energy-conserving nature, leading to inhibition of spontaneous emission Yablonovitch_1987 . In contrast, the non-equilibrium nature of PTCs allows exponentially growing modes in time, enabling light emission and amplification Lyubarov_2022 . This unique property makes temporally engineered band structures highly promising by enabling novel forms of spontaneous emission control Dikopoltsev_2022 ; Park_2025 , overcoming electromagnetic bounds inherent to equilibrium systems Hayran_2024 , and providing non-resonant and tunable gain, lasing, and frequency conversion with reduced phase-matching requirements compared to conventional nonlinear optical methods Lyubarov_2022 ; Khurgin_2024 .
However, experimental realization of a PTCs has also been recognized to be an arduous task Asgari_2024 ; Hayran_2022 ; Saha_2023 ; Hayran_2025 . Much like SPCs require a large dielectric contrast on wavelength scales to reshape the photonic band structure in a sizeable fashion, PTCs demand relative modulation depths of the optical properties to be of the order of unity (e.g. $\delta n/n\approx 1$ with $n$ the refractive index) while simultaneously providing ultrafast coherent modulations at optical cycle rates. From a material perspective Saha_2023 , dielectrics offer ultrafast response times but fail to produce such large modulations depths, while epsilon-near-zero materials produce larger depths but periodically and coherently driven modulations on optical-cycle timescales remains yet to be demonstrated Alam_2016 ; Khurgin_2020 . Furthermore, the estimated $\mathrm{TW/cm^{3}}$ power densities required to produce such nonlinearities are comparable with material damage thresholds, casting doubts on the feasibility Hayran_2022 ; Hayran_2025 . For these reasons, practical implementations of photonic time crystals have so far been ascribed to systems operating at electronic frequencies Reyes_Ayona_2015 ; Wang_2023 ; Lee2025 and the question remains opened as to their possible experimental realization at higher frequencies in an optical context.
@@
Here, we show that this path is experimentally possible and demonstrate the first optical realization of a photonic time crystal. Using a multi-cycle Terahertz-frequency light field of modest amplitude ($\approx 40$kV/cm), we periodically drive a surface plasmonic cavity metamaterial and achieve strong (near-unity) and coherent (sub-optical-cycle) parametric modulation of its resonance, an effect which arises from dynamical variations in the carriers’ effective mass and kinetic energy. Our results provide the basis for time-crystalline phenomena in plasmonic systems Kiselev_2024 ; Feinberg_2025 ; Kiselev2024 , including the parametric amplification of plasmons and the generation of entangled plasmons, all within the experimentally accessible regime of parameters of this platform. Moreover, the combination of metamaterial design and plasmonic tunability Aupiais_2023 ; Aupiais_2024 offers a promising route to enrich PTCs with the full spectrum of photonic degrees of freedom.
@@
To demonstrate our approach, we take a conceptually simple metamaterial structure, namely a periodic lattice of sub-wavelength metal/insulator/plasma cavities, where the plasmonic material corresponds to the small band-gap semiconductor InSb Aupiais_2023 . It is shown in Fig. 1 together with the experimental configuration.
@@
The equilibrium power reflectivity of the metamaterial, shown in Fig. 1b, exhibits an absorption dip centered at $0.77\rm{THz}$ with a full width at half maximum of $\sim 0.18\rm{THz}$. It corresponds to the resonance of the fundamental $\lambda/2$ mode of the sub-wavelength Fabry-Perot surface plasmonic cavity, whose mode profile is displayed in Fig. 1c. The periodically driven regime is investigated using a multi-cycle THz pump pulse centered at $0.69\rm{THz}$—slightly below the equilibrium resonance frequency—with a peak field amplitude of approximately 40kV/cm. The system’s dynamics during the drive is monitored using a broadband THz probe pulse, $\tilde{E}(\tau,t)$, where $\tau$ denotes the delay between the pump and the electro-optic sampling beam, and $t$ represents the internal delay between the sampling beam and the THz probe pulse, as defined in Fig.1a. For each time delay $\tau$, the frequency-resolved response of the metamaterial is obtained by performing a Fourier transform over the internal delay $t$.
@@
We attribute the temporal modulations of the surface plasmon cavity resonance to dynamical variations in the carriers’ effective mass $m^{*}(\tau)$ and kinetic inductance associated with non-parabolic conduction bands in small band-gap semiconductors Houver_2019 ; Soranzio_2024 : $m^{*}(\tau)=m^{*}_{0}+\delta m^{*}(\tau)$, where $m^{*}_{0}$ corresponds to the carriers’ rest effective mass. Under periodic driving at frequency $\omega_{\mathrm{d}}$, the electrons’ momentum $p(\tau)$ oscillates as $p(\tau)=p_{0}\cos(\omega_{\mathrm{d}}\tau)$. For a non-parabolic and isotropic conduction band given by the Kane model Kane_1957 , this results in effective mass modulations at twice the drive frequency:
$\delta m^{*}(\tau)=\alpha p^{2}(\tau)=\frac{\alpha}{2}{p_{0}}^{2}+\frac{\alpha}{2}{p_{0}}^{2}\cos(2\omega_{\mathrm{d}}t)$ (1)
where $\alpha$ accounts for the deviation from a purely parabolic band dispersion ($\alpha>0$ here, see methods). The resulting oscillation of the plasmonic cavity resonance can be qualitatively understood as follows. Using an LC-resonator model for the surface plasmonic cavity, the resonance frequency is given by $\omega_{\mathrm{c}}=\frac{1}{\sqrt{LC}},$ where $L$ and $C$ are the inductance and capacitance of the mode, respectively. For surface plasmons, the inductance $L=L_{G}+L_{K}$ includes contributions from both a geometric inductance $L_{G}$, arising from retardation effects, and a kinetic inductance $L_{K}\propto\frac{m^{*}}{ne^{2}}$, which represents part of the mode energy that is stored into the kinetic energy of the carriers (with $n$ the carrier density and $e$ the charge) Staffaroni_2012 . The kinetic inductance can be separated into an equilibrium and a time-varying part: $L_{K}=L_{K}^{0}+\delta L_{K}(\tau)$. Rigorously, it can be shown that $\delta L_{K}(\tau)/L_{K}^{0}=\beta\delta m^{*}(\tau)/m^{*}_{0}$ where the time-independent parameter $0\leq\beta\leq 1$ reaches unity only in the case of negligible geometric inductance and is lower otherwise (see methods and supplementary material). The resulting modulation of the resonance frequency is:
@@
where the time-dependent modulation parameter is given by $\eta(\tau)=-\frac{\omega_{\rm c}^{0}}{2}\frac{L_{K}^{0}}{L_{G}+L_{K}^{0}}\beta\frac{\delta m^{*}(\tau)}{m^{*}_{0}}$ (see methods and supplementary material). This Hamiltonian describes a parametrically driven cavity resonance and the emergence of an interaction between surface plasmons that is generated by the drive. To capture the metamaterial response function that is measured experimentally -the reflectivity-, we implement an input-output formalism with two baths representing radiative and non-radiative (ohmic) decay channels. These account for coupling to an external electromagnetic probe and material losses, respectively. Adopting the experimental temporal conventions and using a single adjustable parameter, we present in Fig. 2c the computed time-resolved reflection coefficient $r(\tau,\nu)$ that best matches the experimental observations. This approach allows direct comparison between experimental and theoretical response functions, enabling inference of the time-dependent modulation of microscopic system parameters.

### v2

@@ Abstract
Periodically driven optical materials and metamaterials have recently emerged as a promising platform for realizing photonic time crystals (PTCs), which are systems whose optical properties are strongly and periodically modulated on timescales comparable to the optical cycle of light Asgari et al. (2024). These time-varying structures are the temporal counterparts of spatial photonic crystals (SPCs), for which a large and periodic dielectric contrast is achieved spatially on wavelength scales. Just as SPCs have revolutionized control over light–matter interactions by engineering the photonic density of states in space Yablonovitch (1987); John (1987); Joannopoulos et al. (2011); Lu et al. (2014), PTCs promise comparable breakthroughs from a fundamentally new perspective: a temporal one Lustig et al. (2018); Lyubarov et al. (2022); Dikopoltsev et al. (2022); Li et al. (2023); Wang et al. (2024); Park et al. (2025). However, harnessing such phenomena all-optically poses severe experimental challenges Asgari et al. (2024); Khurgin et al. (2020); Hayran et al. (2022); Saha et al. (2023); Hayran and Monticone (2025), as it requires order-unity modulation depths of a material’s optical properties on ultrafast timescales comparable to the light cycle, a regime that has remained elusive to date.
Here, we demonstrate the first all-optical realization of a photonic time crystal, achieved with a surface plasmon cavity metamaterial operating at Terahertz (THz) frequencies. We demonstrate strong (near-unity) and coherent (sub-optical cycle) periodic driving of the plasmonic metamaterial enabled by field-induced dynamical modulation of the carriers’ kinetic energy and effective mass—reaching up to 80% of their rest mass. Our spectroscopic measurements reveal a transition into the PTC regime mediated by an exceptional point, at which two Floquet-driven optical eigenmodes coalesce. In the PTC regime, emergent gain is shown to reduce plasmonic losses by more than 50% Kiselev et al. (2024b); Feinberg et al. (2025) and we predict plasmonic lasing to be within experimental reach. These results pave the way for temporal engineering of losses and light-matter interactions in plasmonic systems, and establish a robust new platform for time-domain photonics.
@@ Introduction
Dynamic control of optical systems in time via periodic modulation, or Floquet driving, has recently emerged as a powerful paradigm to photonic engineering. For a modulation frequency $\omega_{d}$ that is much smaller than the light frequency of interest $\omega$ $(\omega_{d}\ll\omega)$, this approach has been used to engineer a discrete synthetic dimension from the sidebands ladder at $\omega\pm n\omega_{d}$ ($n$ being an integer), a concept formalized within the framework of four-dimensional or Floquet metamaterials Engheta (2020); Yin et al. (2022); Galiffi et al. (2022); Sapienza et al. (2023); Engheta (2023); Horsley and Pendry (2023); Sisler et al. (2024). When the modulation and light frequencies become commensurate together $(\omega_{d}\approx 2\omega)$, qualitatively new phenomena emerge that enable photonic band structures to be engineered dynamically through a temporal approach, an aspect which is captured by the recently introduced concept of photonic time crystals (PTCs) Asgari et al. (2024); Lustig et al. (2018); Lyubarov et al. (2022); Dikopoltsev et al. (2022); Li et al. (2023); Wang et al. (2024); Park et al. (2025); Kiselev et al. (2024b); Feinberg et al. (2025); Pan et al. (2023); Hayran and Monticone (2024); Khurgin (2024). It should be noted that PTCs are periodically driven optical systems that do not break the discrete time-translational invariance imposed by the drive and should thus not be confused with many-body time crystals whose physics originate specifically from the breaking of time-translational invariance Wilczek (2012); Sacha and Zakrzewski (2017). As depicted in Figure 1a, PTCs can be viewed as the temporal analogues of spatial photonic crystals (SPCs) Yablonovitch (1987); John (1987); Joannopoulos et al. (2011); Lu et al. (2014) by permuting space and time coordinates on one hand and momentum and frequency on the other. In SPCs (resp. PTCs), spatial (resp. temporal) periodicity generates band replicas along the momentum (resp. frequency) axis and whose hybridization with the original band opens frequency (resp. momentum) gaps ($\omega$-gaps, resp. $k$-gaps) near their crossings. Wave-behavior within or in the vicinity of these gaps is key in governing light–matter interactions and differ fundamentally between the two cases. Inside the gaps, wave solutions reduce to exponentially decaying or growing modes — spatially in SPCs and temporally in PTCs (Fig. 1a). In SPCs, only the decaying solution is physically allowed, a consequence of their time-invariant and energy-conserving nature, leading to inhibition of spontaneous emission Yablonovitch (1987). In contrast, the non-equilibrium nature of PTCs allows exponentially growing modes in time, enabling light emission and amplification Lyubarov et al. (2022). This unique property makes temporally engineered band structures highly promising for novel forms of spontaneous emission control Dikopoltsev et al. (2022); Park et al. (2025), overcoming electromagnetic bounds inherent to equilibrium systems Hayran and Monticone (2024), and providing non-resonant and tunable gain, lasing, and frequency conversion with reduced phase-matching requirements compared to conventional nonlinear optical methods Lyubarov et al. (2022); Khurgin (2024).
However, all-optical realizations of PTCs have proven to be a challenging task Asgari et al. (2024); Khurgin et al. (2020); Hayran et al. (2022); Saha et al. (2023); Hayran and Monticone (2025). Much like SPCs, which require a large dielectric contrast on wavelength scales to engineer photonic band structures, PTCs demand order-unity relative modulation depths of a material’s optical properties (e.g., $\delta n/n\approx 1$, where $n$ is the refractive index), together with periodic and coherent ultrafast modulations on timescales comparable to the light cycle, two conditions that remain challenging to achieve simultaneously. Consequently, experimental implementations have so far been confined to electrical circuits Reyes-Ayona and Halevi (2015); Wang et al. (2023); Lee et al. (arXiv:2507.19916), leaving open the challenge of achieving PTCs beyond the microwave regime.
@@
Here, we demonstrate the first all-optical realization of a PTC, operating at Terahertz (THz) frequencies and achieved with a surface plasmon cavity metamaterial. Using a multi-cycle THz-frequency light field of modest amplitude ($\approx 40$kV/cm), we periodically drive the plasmonic metamaterial and achieve strong (near-unity) and coherent (sub-optical-cycle) parametric modulation of its resonance, an effect which arises from dynamical variations in the carriers’ effective mass and reaching up to 80% of their rest mass. Via spectroscopic measurements, we uncover the transition to the PTC regime as a function of time which manifests, in particular, as a substantial decrease of the plasmonic losses by more than 50%, opening the door to the temporal engineering of plasmonic losses and plasmonic lasing in this system Kiselev et al. (2024b); Feinberg et al. (2025); Kiselev et al. (2024a).
@@
To demonstrate our approach, we take a conceptually simple metamaterial structure, namely a periodic lattice of sub-wavelength metal/insulator/plasma cavities, where the plasmonic material corresponds to the small band-gap semiconductor InSb Aupiais et al. (2023). It is shown in Fig. 1b together with the experimental configuration.
@@
The equilibrium power reflectivity of the metamaterial, shown in Fig. 1c, exhibits an absorption dip centered at $0.77\rm{THz}$ with a full width at half maximum of $\sim 0.18\rm{THz}$. It corresponds to the resonance of the fundamental $\lambda/2$ mode of the sub-wavelength Fabry-Perot surface plasmonic cavity, whose mode profile is displayed in Fig. 1d. The periodically driven regime is investigated using a narrowband THz pump pulse centered at $0.69\rm{THz}$—slightly below the equilibrium resonance frequency—with a peak field amplitude of approximately 40kV/cm. The system’s dynamics during the drive is monitored by a broadband THz probe pulse, $\tilde{E}(\tau,t)$, where $\tau$ denotes the delay between the pump and the electro-optic sampling beam, and $t$ represents the internal delay between the sampling beam and the THz probe pulse, as defined in Fig. 1b. For each time delay $\tau$, the frequency-resolved response of the metamaterial is obtained by performing a Fourier transform over the internal delay $t$.
@@
We attribute the temporal modulations of the surface plasmon cavity resonance to dynamical variations in its kinetic inductance through the carriers’ effective mass $m^{*}(\tau)$, which we write $m^{*}(\tau)=m_{0}^{*}+\delta m^{*}(\tau)$, where $m_{0}^{*}$ denotes the carriers’ rest effective mass. Such variations are known to arise from non-parabolic conduction bands in narrow band-gap semiconductors Houver et al. (2019); Soranzio et al. (2024) and have also been discussed in the context epsilon-near-zero plasmonic materials operating at optical frequencies Khurgin et al. (2020).
Under periodic driving at frequency $\omega_{\mathrm{d}}$, the electrons’ momentum $p(\tau)$ oscillates as $p(\tau)=p_{0}\cos(\omega_{\mathrm{d}}\tau)$. For a non-parabolic and isotropic conduction band given by the Kane model Kane (1957), this results in effective mass modulations at twice the drive frequency:
$\delta m^{*}(\tau)=\alpha p^{2}(\tau)=\frac{\alpha}{2}{p_{0}}^{2}+\frac{\alpha}{2}{p_{0}}^{2}\cos(2\omega_{\mathrm{d}}\tau)$ (1)
where $\alpha$ accounts for the deviation from a purely parabolic band dispersion ($\alpha>0$ here, see methods). The resulting oscillation of the plasmonic cavity resonance can be qualitatively understood as follows. Using an LC-resonator model for the surface plasmonic cavity, the resonance frequency is given by $\omega_{\mathrm{c}}=\frac{1}{\sqrt{LC}},$ where $L$ and $C$ are the inductance and capacitance of the mode, respectively. For surface plasmons, the inductance $L=L_{G}+L_{K}$ includes contributions from both a geometric inductance $L_{G}$, arising from retardation effects, and a kinetic inductance $L_{K}\propto\frac{m^{*}}{ne^{2}}$, which represents part of the mode energy that is stored into the kinetic energy of the carriers (with $n$ the carrier density and $e$ the charge) Staffaroni et al. (2012). The kinetic inductance can be separated into an equilibrium and a time-varying part: $L_{K}=L_{K}^{0}+\delta L_{K}(\tau)$. Rigorously, it can be shown that $\delta L_{K}(\tau)/L_{K}^{0}=\beta\delta m^{*}(\tau)/m^{*}_{0}$ where the time-independent parameter $0\leq\beta\leq 1$ reaches unity only in the case of negligible geometric inductance and is lower otherwise (see methods and supplementary material). The resulting modulation of the resonance frequency is:
@@
where the time-dependent modulation parameter is given by $\eta(\tau)=-\frac{\omega_{\rm c}^{0}}{2}\frac{L_{K}^{0}}{L_{G}+L_{K}^{0}}\beta\frac{\delta m^{*}(\tau)}{m^{*}_{0}}$ (see methods and supplementary material). To capture the metamaterial response function that is measured experimentally -the reflectivity-, we implement an input-output formalism with two baths representing radiative and non-radiative (ohmic) decay channels. These account for coupling to an external electromagnetic probe and material losses, respectively. Adopting the experimental temporal conventions and using a single adjustable parameter, we present in Fig. 2c the computed time-resolved reflection coefficient $\tilde{r}(\tau,\omega)$ that best matches the experimental observations. This approach allows direct comparison between experimental and theoretical response functions, enabling inference of the time-dependent modulation of microscopic system parameters.
Figure 3: a) Real and imaginary parts of the zero-order (center panel) and first-orders (outer panels) Floquet spectra obtained from Fourier Transform of the time domain data of Fig 2b. over one drive period at $\tau\approx 24ps$. Experimental data points are shown together with a fitting of the lineshapes (see text). b) Evolution of the complex-frequency poles of the reflectivity, controlled by the Floquet eigenmodes, as a function of drive strength across the PTC transition. c) Real and imaginary parts of $\Delta\tilde{\mu}/\omega_{c}^{0}$ as a function of drive strength $\eta_{0}/\omega_{c}$ computed for parameters taken from the experiment: $\omega_{d}/\omega_{c}^{0}=0.92$, $\Gamma/\omega_{c}^{0}=0.12$. Dashed lines indicate the values of $\eta_{0}/\omega_{c}$ taken for the computation of the spectra shown in Fig. 3d. d) Spectra of the real part of $\tilde{r}_{0}(\omega)$ and imaginary part of $\tilde{r}_{-1}(\omega)$ across the PTC transition and computed for values of $\eta_{0}/\omega_{c}$ indicated by the dashed lines in panel c.